# Sigma model renormalisation group flows, singularities and some remarks on cosmology

### v1

@@ 1 Introduction
$L=\frac{1}{2}\,g(\phi)_{\mu\nu}\,\eta^{ab}\partial_{a}\phi^{\mu}\,\partial_{b}\phi^{\nu}\leavevmode\nobreak\ ,$ (1.1)
@@
The question that will be explored is the following: what is the behaviour of RG flows as either $\mu\rightarrow 0$ (equivalently $t\rightarrow+\infty)$) or $\mu\rightarrow+\infty$ (equivalently $t\rightarrow$-$\infty$)?. The emphasis will be on the former, i.e. what does it happen to the metric $g$ as the energy scale decreases starting from a generic metric $g$ on the sigma model manifold $M^{n}$? In particular, suppose that one fixes the topology of $M^{n}$ and starts at a scale $\mu_{0}$ with a generic metric $g_{0}$, how does one expect that $g_{0}$ will evolve with (1.2) as the energy scale goes into infrared? The analogous question with the scale going into UV can be analysed in a similar way. This also applies to manifolds $M^{n}$ that potentially admit a high degree of symmetry, like the spheres, but the starting point can be a metric $g_{0}$ that may not admit any isometries – the behaviour of the RG flow depends on the initial metric $g_{0}$.
@@
It is widely accepted that the universe has a small positive, but not vanishing, cosmological constant, $\Lambda\sim 10^{-123}l^{-2}_{\mathrm{Pl}}$, that drives its expansion and there are recent suggestions that this may have not been constant throughout the lifetime of the universe. The difficulty of incorporating such a small cosmological constant in field theory and a cosmological constant in string theory, is well documented, for string theory see the proposals of e.g. [25] and [26]. Here, making the hypothesis that the universe satisfies an RG flow equation as in (1.2), we shall demonstrate that spacetimes with a positive cosmological constant can be incorporated in a string theory setting. We shall achieved this by utilising the properties of the Ricci solitons that are solutions to the sigma model RG flow. In particular, we shall show that de-Sitter space is such a solution. Moreover, as Ricci solitons can model the geometry of the spacetime near certain RG flow singularities and these singularities can occur at different regions of the spacetime, we shall hypothesise that potentially different cosmological constants can occur at different epochs.
@@
The standard cosmological model is based on the observation that the universe at very large scales is homogeneous and isotropic. The mathematical models for the geometry of the universe have built in this symmetry from the beginning all the way back to the very origin of the universe. However, this need not be the case. It is not clear why from the vast number of geometries that the universe may have had as it emerged from the quantum epoch, presumably described by a theory of quantum gravity, to the geometric era, described by general relativity, only those that exhibit such a high degree of symmetry have occurred. Therefore, it would be useful if there is a mechanism, even if it is for illustration purpose only, i.e. toy model, that could allow from the vast number of possibilities that can occur as geometries for the universe to select the ones that exhibit a high degree of symmetry. Of course, this is reminiscent to the question answered in the Perelman’s proof [6] of Thurston’s geometrisation conjecture [27]. Thus, the question that we raise is whether the sigma model RG flow together with the work of Thurston, Hamilton and Perelman can serve to provide such a proof for the homogeneity and isotropy of the universe.
@@ 2.1.1 Ricci solitons and singularities
$R_{\mu\nu}=\Lambda g_{\mu\nu}+\nabla_{\mu}X_{\nu}+\nabla_{\nu}X_{\mu}\leavevmode\nobreak\ ,$ (2.1)
where $\Lambda$ is a constant [^2] We shall often refer to $\Lambda$ as the cosmological constant. However, in $n>2$ dimensions the cosmological constant, $\Lambda_{c}$, as it appears in the Einstein equations, $G+\Lambda_{c}g=\kappa T$, is related to $\Lambda$ as $\Lambda=\frac{2\Lambda_{c}}{n-2}$, where $G$ is the Einstein tensor. and $X=X^{\mu}\partial_{\mu}$ is a vector field on $M^{n}$ with $X_{\mu}=g_{\mu\nu}X^{\nu}$. The solitons with $\Lambda<0$, $\Lambda=0$ and $\Lambda>0$ are expanding, steady and shrinking, respectively. Moreover, if $X=-d\Phi$, then the Ricci solitons are called gradient soliton. As we shall mention below, type I singularities in the Ricci flow are related to Ricci solitons – the latter provide a model for the geometry of the manifold $M^{n}$ near the type I singularities.
@@
$\Omega^{2}(t)=1-\Lambda t\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ Y_{t}(x)=-\Omega^{-2}X(x)\leavevmode\nobreak\ .$ (2.3)
@@
$R_{\mu\nu}+2\nabla_{\mu}\partial_{\nu}\Phi=0\leavevmode\nobreak\ ,$ (2.4)
@@
$g(t)_{\mu\nu}=-\big{(}\Lambda t-1\big{)}g_{\mu\nu}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\$(2.6)
@@
$g=du^{2}+\tanh^{2}u\,d\theta^{2}\leavevmode\nobreak\ .$ (2.8)
@@
$X=-\tanh u\,\partial_{u}\leavevmode\nobreak\ .$ (2.10)
@@
$\sinh u(t)=e^{-t}\sinh u_{0}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \theta=\theta_{0}\leavevmode\nobreak\ ,$ (2.11)
@@
$g(t)=f_{t}^{*}g_{0}=du^{2}+\tanh^{2}u(t)\,d\theta^{2}=\frac{e^{-2t}\cosh^{2}u_{0}}{1+e^{-2t}\sinh^{2}u_{0}}\Big{(}du_{0}^{2}+\tanh^{2}u_{0}\,d\theta_{0}^{2}\Big{)}\leavevmode\nobreak\ .$ (2.12)
@@ 2.1.2 Type I singularities
$\sup_{M^{n}\times[0,T)}\big{(}(T-t)|R(x,t)|\big{)}\leq C<\infty\leavevmode\nobreak\ ,$ (2.13)
@@
$\sup_{M^{n}\times(T,0]}\big{(}(t-T)|R(x,t)|\big{)}\leq C<\infty\leavevmode\nobreak\ ,$ (2.19)
@@ 2.2 Application to sigma model RG flows
The results we have described for the behaviour of the solutions to the Ricci flow can be easily adapted to describe the behaviour of RG flow in two-dimensional sigma models. As it has already be mentioned the Ricci flow parameter $t$ is related to the energy scale $\mu$ of the two-dimensional sigma models with $t=-\log\mu$. From the string perspective, the collective [^8] Sigma models have infinite many couplings given by the components of the spacetime metric at every point of the spacetime or equivalently sigma model manifold $M^{n}$. The ratio $\ell_{s}/R_{M}$ provides an overall description of the couplings of the sigma model as the two-dimensional energy scale changes. sigma model coupling is the ratio $\ell_{s}/R_{M}$ of string length $\ell_{s}$ with the radius $R_{M}$ of the sigma model manifold $M^{n}$. Therefore, at strong sigma model coupling $\ell_{s}>>R_{M}$. Independently on whether this happens at the UV or IR, the (length of the) curvature of the spacetime blows up and so the description of the string effective theory [^9] Here, we assume an off-shell formulation of string theory away from the usual description that the sigma model is taken to be a conformal field theory, for a review see [31]. breaks down.
@@
$g_{i\mu\nu}(s,x)=\tilde{Q}_{i}\,g_{\mu\nu}(t_{i}+\frac{s}{\tilde{Q}_{i}},x)\leavevmode\nobreak\ ,$ (2.20)
@@ 3.1 Cosmological constants
$R(g_{\mathrm{d}S}(t))_{\mu\nu}=R(g_{\mathrm{d}S})_{\mu\nu}=\Lambda(g_{\mathrm{d}S})_{\mu\nu}=\frac{\Lambda}{1-\Lambda t}(g_{\mathrm{d}S}(t))_{\mu\nu}\leavevmode\nobreak\ .$ (3.1)
@@
$\frac{dg(t)_{\mu\nu}}{dt}$ $=$ $-\Big{(}R(g(t))_{\mu\nu}-\kappa\tilde{T}(\phi(t),g(t))_{\mu\nu}\Big{)}\leavevmode\nobreak\ ,$ (3.2)
$\frac{d\phi(t)}{dt}$ $=$ $-{\mathcal{F}}(\phi(t),\partial\phi(t),\partial^{2}\phi(t),g(t))\leavevmode\nobreak\ ,$ (3.3)
@@
In the context of string theory, the flow parameter $t$ is related to the energy scale $\mu$ of the two-dimensional sigma model that describes the propagation of a string in a spacetime. This persists after the inclusion of other string theory fields like the $B$-field as sigma model couplings. However, if the flow equations (3.3) do not have a string theory origin, the question will arise about the interpretation of the flow parameter $t$. There is no straightforward answer, we know, to this question. Though, it may be possible to relate $t$ to thermodynamic quantities that arise in the description of the universe in cosmology like temperature or energy density.
@@ 3.2 Homogeneity and isotropy of the universe
The homogeneity and isotropy of the universe at large scales is built in the cosmological models from the very beginning. However, one can raise the question whether the universe emerged from its quantum epoch with such a high degree of symmetry. It is likely that the geometry of the universe in the beginning of its geometric phase had regions with positive and negative curvature and not exhibiting any particular symmetry. This is a reasonable assumption as otherwise it would have involved fine tuning to select such a highly symmetric geometry from the vast number of other possibilities. If we accept this, then the question arises how the universe evolved to the current symmetric state we observe at large scales. This question is reminiscent of the question posed in Thurston’s geometrisation conjecture for 3-manifolds proven by Perelman. This classifies the “simplest” geometries that can be put on (compact without boundary oriented) 3-dimensional manifolds allowed by their topology as that described by their fundamental group. There are 8 such “simple” geometries in three dimensions, all admitting a transitive group action, which can be combined in a certain way to produce a “simple” geometry on any 3-dimensional manifold. Of course, all 3-dimensional manifolds can admit metrics that do not admit any symmetries.
@@
$\frac{d}{dt}\gamma_{ij}$ $=$ $-\Big{(}R_{ij}^{(3)}+KK_{ij}-2K_{ip}K^{p}{}_{j}+\frac{1}{N}(\partial_{\tau}K_{ij}-\mathcal{L}_{N}K_{ij}+D_{i}D_{j}N)\Big{)}\leavevmode\nobreak\ ,$ (3.5)
@@
$K_{ij}=\frac{1}{2N}(\partial_{\tau}\gamma_{ij}-D_{i}N_{j}-D_{j}N_{i})\leavevmode\nobreak\ ,$ (3.6)
@@ Acknowledgements
I thank Andreas Stergiou for helpful discussions.

### v2

@@ 1 Introduction
$L=\frac{1}{2}\,g(\phi)_{\mu\nu}\,\eta^{ab}\partial_{a}\phi^{\mu}\,\partial_{b}\phi^{\nu}\penalty 10000\ ,$ (1.1)
@@
The question that will be explored is the following: what is the behaviour of RG flows as either $\mu\rightarrow 0$ (equivalently $t\rightarrow+\infty)$) or $\mu\rightarrow+\infty$ (equivalently $t\rightarrow$-$\infty$)? The emphasis will be on the former, i.e. what does it happen to the metric $g$ as the energy scale decreases starting from a generic metric $g$ on the sigma model manifold $M^{n}$? In particular, suppose that one fixes the topology of $M^{n}$ and starts at a scale $\mu_{0}$ with a generic metric $g_{0}$, how does one expect that $g_{0}$ will evolve with (1.2) as the energy scale goes into infrared? The analogous question with the scale going into UV can be analysed in a similar way. This also applies to manifolds $M^{n}$ that potentially admit a high degree of symmetry, like the spheres, but the starting point can be a metric $g_{0}$ that may not admit any isometries – the behaviour of the RG flow depends on the initial metric $g_{0}$.
@@
It is widely accepted that the universe has a small positive, but not vanishing, cosmological constant, $\Lambda\sim 10^{-123}l^{-2}_{\mathrm{Pl}}$, that drives its expansion. The difficulty of incorporating such a small cosmological constant in field theory and a cosmological constant in string theory, is well documented, for string theory see the proposals of e.g. [25] and [26]. Here, making the hypothesis that the universe satisfies an RG flow equation as in (1.2), we shall demonstrate that spacetimes with a positive cosmological constant can be incorporated in a string theory setting. We shall achieve this by utilising the properties of the Ricci solitons that are solutions to the sigma model RG flow. In particular, we shall show that de-Sitter space is such a solution. Moreover, as Ricci solitons can model the geometry of the spacetime near certain RG flow singularities and these singularities can occur at different regions of the spacetime, we shall hypothesise that potentially different cosmological constants can occur at different epochs. The use of renormalisation group flow to model properties of spacetime has arisen before in the context [^2] There are some apparent differences in the use of RG proposed in this paper and that in AdS/CFT and holographic cosmology. The RG flow in AdS/CFT and holographic cosmology is that of the dual gauge theory while in this paper is the Ricci flow of the metric or equivalently the RG flow of the sigma model. Though, there may still be a relation to be explored in the future. . In the former, the bulk coordinate transverse to those of the boundary that the gauge theory is localised on has been interpreted as the energy scale of the gauge theory [27], while in the latter the energy scale of the holographic gauge theory is identified with the inverse of the radius of the universe [28], see also [29] for a concise exposition.
@@
The standard cosmological model is based on the observation that the universe at very large scales is homogeneous and isotropic, which is also known as the horizon problem. The accepted explanation for this is the inflationary scenario for the universe [30, 31], where it is proposed that all the observable universe today is the result of the exponential expansion of a small region in the early universe that was in causal contact – this explains why the temperature of the microwave radiation is almost uniform in every direction. The mathematical models for the geometry of the universe have built in this symmetry from the beginning all the way back to the very origin of the universe – this includes some calculations within the inflationary scenario. However, this need not be the case. It is not clear why from the vast number of geometries that the universe may have had as it emerged from the quantum epoch, presumably described by a theory of quantum gravity, to the geometric era, described by (semi-classical) general relativity, only those that exhibit such a high degree of symmetry have occurred. Therefore, it would be useful if there is a mechanism, even if it is for illustration purpose only, i.e. toy model, that could allow from the vast number of possibilities that can occur as geometries for the universe to select the ones that exhibit a high degree of symmetry. Of course, this is reminiscent to the question answered in the Perelman’s proof [6] of Thurston’s geometrisation conjecture [32]. Thus, the question that we raise is whether the sigma model RG flow together with the work of Thurston, Hamilton and Perelman can serve to provide such a proof for the homogeneity and isotropy of the universe.
@@ 2.1.1 Ricci solitons and singularities
$R_{\mu\nu}=\Lambda g_{\mu\nu}+\nabla_{\mu}X_{\nu}+\nabla_{\nu}X_{\mu}\penalty 10000\ ,$ (2.1)
where $\Lambda$ is a constant [^3] We shall often refer to $\Lambda$ as the cosmological constant. However, in $n>2$ dimensions the cosmological constant, $\Lambda_{c}$, as it appears in the Einstein equations, $G+\Lambda_{c}g=\kappa T$, is related to $\Lambda$ as $\Lambda=\frac{2\Lambda_{c}}{n-2}$, where $G$ is the Einstein tensor. and $X=X^{\mu}\partial_{\mu}$ is a vector field on $M^{n}$ with $X_{\mu}=g_{\mu\nu}X^{\nu}$. The solitons with $\Lambda<0$, $\Lambda=0$ and $\Lambda>0$ are expanding, steady and shrinking, respectively. Moreover, if $X=-d\Phi$, then the Ricci solitons are called gradient solitons. As we shall mention below, type I singularities in the Ricci flow are related to Ricci solitons – the latter provide a model for the geometry of the manifold $M^{n}$ near the type I singularities.
@@
$\Omega^{2}(t)=1-\Lambda t\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ Y_{t}(x)=-\Omega^{-2}X(x)\penalty 10000\ .$ (2.3)
@@
$R_{\mu\nu}+2\nabla_{\mu}\partial_{\nu}\Phi=0\penalty 10000\ ,$ (2.4)
@@
$g(t)_{\mu\nu}=-\big(\Lambda t-1\big)g_{\mu\nu}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\$(2.6)
@@
$g=du^{2}+\tanh^{2}u\,d\theta^{2}\penalty 10000\ .$ (2.8)
@@
$X=-\tanh u\,\partial_{u}\penalty 10000\ .$ (2.10)
@@
$\sinh u(t)=e^{-t}\sinh u_{0}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \theta=\theta_{0}\penalty 10000\ ,$ (2.11)
@@
$g(t)=f_{t}^{*}g_{0}=du^{2}+\tanh^{2}u(t)\,d\theta^{2}=\frac{e^{-2t}\cosh^{2}u_{0}}{1+e^{-2t}\sinh^{2}u_{0}}\Big(du_{0}^{2}+\tanh^{2}u_{0}\,d\theta_{0}^{2}\Big)\penalty 10000\ .$ (2.12)
@@ 2.1.2 Type I singularities
$\sup_{M^{n}\times[0,T)}\big((T-t)|R(x,t)|\big)\leq C<\infty\penalty 10000\ ,$ (2.13)
@@
$\sup_{M^{n}\times(T,0]}\big((t-T)|R(x,t)|\big)\leq C<\infty\penalty 10000\ ,$ (2.19)
@@ 2.2 Application to sigma model RG flows
The results we have described for the behaviour of the solutions to the Ricci flow can be easily adapted to describe the behaviour of RG flow in two-dimensional sigma models. As it has already been mentioned the Ricci flow parameter $t$ is related to the energy scale $\mu$ of the two-dimensional sigma models with $t=-\log\mu$. From the string perspective, the collective [^9] Sigma models have infinite many couplings given by the components of the spacetime metric at every point of the spacetime or equivalently sigma model manifold $M^{n}$. The ratio $\ell_{s}/R_{M}$ provides an overall description of the couplings of the sigma model as the two-dimensional energy scale changes. sigma model coupling is the ratio $\ell_{s}/R_{M}$ of string length $\ell_{s}$ with the radius $R_{M}$ of the sigma model manifold $M^{n}$. Therefore, at strong sigma model coupling $\ell_{s}>>R_{M}$. Independently on whether this happens at the UV or IR, the (length of the) curvature of the spacetime blows up and so the description of the string effective theory [^10] Here, we assume an off-shell formulation of string theory away from the usual description that the sigma model is taken to be a conformal field theory, for a review see [36]. breaks down.
@@
$g_{i\mu\nu}(s,x)=\tilde{Q}_{i}\,g_{\mu\nu}(t_{i}+\frac{s}{\tilde{Q}_{i}},x)\penalty 10000\ ,$ (2.20)
@@ 3.1 Cosmological constants
$R(g_{\mathrm{d}S}(t))_{\mu\nu}=R(g_{\mathrm{d}S})_{\mu\nu}=\Lambda\,(g_{\mathrm{d}S})_{\mu\nu}=\frac{\Lambda}{1-\Lambda t}(g_{\mathrm{d}S}(t))_{\mu\nu}\penalty 10000\ .$ (3.1)
@@
$\frac{dg(t)_{\mu\nu}}{dt}$ $=$ $-\Big(R(g(t))_{\mu\nu}-\kappa\tilde{T}(\phi(t),g(t))_{\mu\nu}\Big)\penalty 10000\ ,$ (3.2)
$\frac{d\phi(t)}{dt}$ $=$ $-{\mathcal{F}}(\phi(t),\partial\phi(t),\partial^{2}\phi(t),g(t))\penalty 10000\ ,$ (3.3)
@@
In the context of string theory, the flow parameter $t$ is related to the energy scale $\mu$ of the two-dimensional sigma model that describes the propagation of a string in a spacetime. This persists after the inclusion of other string theory fields like the $B$-field as sigma model couplings. However, if the flow equations (3.3) do not have a string theory origin, the question will arise about the interpretation of the flow parameter $t$. There is no straightforward answer, we know, to this question. However, on consideration of holographic cosmology [37, 38, 39, 40], where the quantum theory of gravity has been explored with asymptotic de Sitter boundary conditions, and in view of the results of [27] and [28], where the energy scale of the dual field theory is identified the transverse coordinate of the bulk, t-t could be identified with cosmological time $\tau$. However, it should be emphasised that this is an analogy. Nevertheless, such an identification has implications that require further investigation to be reported elsewhere.
@@ 3.2 Homogeneity and isotropy of the universe
The homogeneity and isotropy of the universe at large scales is built in the cosmological models from the very beginning. However, one can raise the question whether the universe emerged from its quantum epoch with such a high degree of symmetry. It is likely that the geometry of the universe in the beginning of its geometric phase had regions with positive and negative curvature and not exhibiting any particular symmetry. This is a reasonable assumption as otherwise it would have involved fine tuning to select such a highly symmetric geometry from the vast number of other possibilities [^15] In the inflationary universe, this issue is resolved because our current observable universe originates from the exponential expansion of a small patch in the past that was in causal contact leading to thermal equilibrium. This is used to explain the uniform temperature for the microwave radiation we see in all directions. However here, we take a geometric approach to the problem.. If we accept this, then the question arises how the universe evolved to the current symmetric state we observe at large scales. This question is reminiscent of the question posed in Thurston’s geometrisation conjecture for 3-manifolds proven by Perelman. This classifies the “simplest” geometries that can be put on (compact without boundary oriented) 3-dimensional manifolds allowed by their topology as that described by their fundamental group. There are 8 such “simple” geometries in three dimensions, all admitting a transitive group action, which can be combined in a certain way to produce a “simple” geometry on any 3-dimensional manifold. Of course, all 3-dimensional manifolds can admit metrics that do not admit any symmetries.
@@
$\frac{d}{dt}\gamma_{ij}$ $=$ $-\Big(R_{ij}^{(3)}+KK_{ij}-2K_{ip}K^{p}{}_{j}+\frac{1}{N}(\partial_{\tau}K_{ij}-\mathcal{L}_{N}K_{ij}+D_{i}D_{j}N)\Big)\penalty 10000\ ,$ (3.5)
@@
$K_{ij}=\frac{1}{2N}(\partial_{\tau}\gamma_{ij}-D_{i}N_{j}-D_{j}N_{i})\penalty 10000\ ,$ (3.6)
@@ Acknowledgements
I thank Kostas Skenderis and Andreas Stergiou for helpful discussions.